Can someone help me solve this?
Type of Problem: Applications of Linear Systems
Solve the problem by using a system of equations.
The school student council voted to go to a baseball game for their annual trip. They bought 30 tickets. Some of the tickets cost $21 each, and some cost $27 each. The total cost of all the tickets was $696. How many of the $21 tickets did they buy?

Answer by rothauserc(4718)

x + y = 30
x = (30-y)
21*(30-y) +27y = 696
630 -21y +27y = 696
6y = 66
y = 11
x = 30 -11 = 19
19, $21 tickets were bought